ganeshie8 (ganeshie8):

prove \[\lim_{t\to0}\dfrac{f(x)^t-1}{t}=\ln(f(x))\]

ganeshie8 (ganeshie8):

@eliassaab

OpenStudy (kainui):

Looks suspiciously like: \[\Large \frac{y^t-y^0}{t-0}\]

ganeshie8 (ganeshie8):

Ahh limit definition of derivative ? nice nice let me try

ganeshie8 (ganeshie8):

\[[f(x)^t ]'(0) = f(x)^t \ln f(x) (0) = \ln f(x)\] \(\blacksquare\) ?

OpenStudy (kainui):

\[\frac{\partial}{\partial y}(f(x)^y)=\lim_{t \to 0} \frac{f(x)^{y+t}-f(x)^y}{t} \\f(x)^y \lim_{t \to 0} \frac{f(x)^t-1}{t} = f(x)^y \ln(f(x)) \\ \lim_{t \to 0} \frac{f(x)^t-1}{t} = \ln(f(x))\]

OpenStudy (kainui):

OS lags too much for me these days to use it fast enough lol. XD I keep having to refresh to use the Post button.

hartnn (hartnn):

since t is the variable, f(x) is a constant, say 'a' so this gets converted into a standard limit.... which can be proved even by L'Hopital

ganeshie8 (ganeshie8):

same here :| the replies are dancing on my end lol anyway to work this without LH or derivatives ? cuz in the previous post we wanted to work it without LH

OpenStudy (anonymous):

Does this "prove" the limit though?

ganeshie8 (ganeshie8):

idk it looks like a legit proof to me if we believe in derivatives ?

OpenStudy (kainui):

I didn't really explain my version fairly well since I skipped steps that seem crucial. I am basically saying I'm using the chain rule for the right hand side of the second line of my "proof" since that's all it really is.

ganeshie8 (ganeshie8):

@SithsAndGiggles we wanted to work this without LH/derivatives in the previous post, i am still trying to cook up a proof using real analysis stuff like epsilon/delta, cauchy or opensets.. but no significant progress as such yet

OpenStudy (anonymous):

This limit does not depend on x, so you can call a=f(x). Let h(t) = a^t, then h'(t) = a^t ln(a) the limit is h'(0)= ln(a)=ln(f(x))

OpenStudy (anonymous):

I'm not sure about this myself, but does the squeeze theorem "prove" a limit exists, or is it just a way of computing a limit?

OpenStudy (anonymous):

Wolframalpha knows that too http://www.wolframalpha.com/input/?i=limit+%28+f%28x%29%5Et+-1%29%2Ft+when+t+goes+to+zero

OpenStudy (anonymous):

You are just reproving some known fact http://www.wyzant.com/resources/lessons/math/calculus/derivative_proofs/a_to_the_x

ganeshie8 (ganeshie8):

I see other ways to work this only complicate things ;) thanks everyone !

OpenStudy (anonymous):

nice nice , was late to this lol

ganeshie8 (ganeshie8):

gimme an analysis proof after your exams @Marki

OpenStudy (anonymous):

u wanna question , or wanna me to answer them B-) ?

OpenStudy (anonymous):

or you wanna how to prove this with analysis

ganeshie8 (ganeshie8):

look at robjohn's reply

OpenStudy (anonymous):

i wonder what condition on f(x) , does it work when f(something )=0 it should be f(x)>0 , right ?

OpenStudy (anonymous):

Um, what in the world is this? You're not making no sense.

OpenStudy (anonymous):

oh so u have other sense about this :P *doubt*